Bonus: notice how we have cleverly used the comb of bounce frequencies around the calibration line to determine that REFL11 is clipping!
Rana and I noticed last week that it looked like the REFL11 beam was clipping. This afternoon, I locked the PRMI with REFL 165 I&Q, and checked the REFL 11 path. The beam looks fine through all of the optics going to the diode, so I just realigned the beam onto the diode using the itty bitty steering mirror. I have not yet checked the change (hopefully improvement) in the REFL11 spectrum.